# Towards Scalable Automated Alignment of LLMs: A Survey

### v1

@@ 1 Introduction
Recent years have witnessed the rapid advancements of large language models (LLMs), which have dramatically reshaped the landscape of artificial intelligence (Ouyang et al., 2022; Touvron et al., 2023; OpenAI, 2023b). Alignment is at the core of shaping behaviors of LLMs corresponding to human intentions and values (Yao et al., 2023a; Shen et al., 2023b), e.g., teaching LLMs to follow “helpful, harmless and honest (HHH)” principles during responding (Askell et al., 2021). As a result, increasing efforts have been made for aligning LLMs to meet the human requirements, which makes it a hotspot research direction in LLM era (Wang et al., 2023g, 2024g; Ji et al., 2023).
@@
To address these challenges, automated alignment has drawn great attention very recently (Yuan et al., 2024; Chen et al., 2024g). Unlike previous methods that relied on human annotation to obtain alignment signals, the goal of automated alignment is constructing scalable and high-quality alignment systems with minimal human intervention. Therefore, automated alignment has the potential to address the core challenges posed by the rapid development of LLMs, where human annotation is either infeasible or extremely expensive. For automated alignment, the most crucial part is to find a scalable alignment signal that can replace human manually-created preference signals and remain effective amid the rapid development of LLMs.
@@
Bias §3, for tree=acquisition [From features of LLMs, [Uncertainty Filtering, acquisition [Self-Consistency (Wang et al., 2023e); Self-Improve (Huang et al., 2023a); West-of-N (Pace et al., 2024); etc., acquisition_work] ] [Self-Judge/Critic/Refine, acquisition [ Constitutional AI (Bai et al., 2022b); Tree-of-Thought (Yao et al., 2023b); Self-Rewarding (Yuan et al., 2024); etc.,acquisition_work ] ] [Context Distillation, acquisition [Lab for Alignment (Askell et al., 2021); Dromedary (Sun et al., 2023d); Llama-2-Chat (Touvron et al., 2023); RLCD (Yang et al., 2024b); etc., acquisition_work] ] ] [From organization of LLMs, [Task Decomposition, acquisition [Least-to-Most (Zhou et al., 2023b); IDA (Christiano et al., 2018); etc., acquisition_work] ] [Self-play, acquisition [SPIN (Chen et al., 2024g); Consensus Game (Jacob et al., 2024); etc., acquisition_work] [Debate (Irving et al., 2018); SPAG (Cheng et al., 2024); etc., acquisition_work] ] ] ] [Behavior
Imitation §4, for tree=representation [ Instruction Construction, [ Unnatural Instructions (Honovich et al., 2023); Self-Instruct (Wang et al., 2023f); Evol-Instruct (Xu et al., 2024a); Humpback (Li et al., 2024e); etc., representation_work_2 ] ] [ Strong-to-Weak Distillation, [Response-guided, representation [ LLaMA-GPT4 (Peng et al., 2023); Stanford Alpaca (Taori et al., 2023); Ultrachat (Ding et al., 2023); etc., representation_work ] ] [Preference-guided, representation [ Zephyr (Tunstall et al., 2023); IterAlign (Chen et al., 2024e); Openchat (Wang et al., 2024b); etc., representation_work ] ] ] [ Weak-to-Strong Alignment, [ Weak2Strong (Burns et al., 2023); IaR (Somerstep et al., 2024); Liu and Alahi (2024); Hase et al. (2024); etc., representation_work_2 ] ] ] [Model
@@
Feedback §6, for tree=editing [ Social Interactions, editing [StableAlignment (Liu et al., 2023a); MoralDial (Sun et al., 2023a); SOTOPIA-$\pi$ (Wang et al., 2024f); etc., editing_work] ] [Human
Collective Reason, editing [Constitutional AI (Bai et al., 2022b); Collective Constitutional AI (Anthropic, 2023) etc., editing_work] ] [Tool Execution, editing [Self-Debugging (Qiao et al., 2024); CodeRL (Qiao et al., 2024); SelfEvlove (Jiang et al., 2023); CRITIC (Gou et al., 2024); etc., editing_work ] ] [Embodied Environment, editing [GLAM (Carta et al., 2023); E2WM (Xiang et al., 2023); TWOSOME (Tan et al., 2024); etc., editing_work ] ] ] [Mechanism §7, for tree=application [Alignment
Mechanism, application [LIMA (Zhou et al., 2023a); Rethinking (Ren et al., 2024); URIAL (Lin et al., 2024a); ICL&IT (Duan et al., 2023); Behavior shift (Wu et al., 2023a) etc., application_work] ] [Inner Workings of Self-feedback,application [GV-consistency (Li et al., 2024f); CriticBench (Lin et al., 2024b); Self-Rewarding (Yuan et al., 2024); Humback (Li et al., 2024e); LLM-as-a-Judge (Zheng et al., 2024b) etc., application_work] ] [Feasibility of Weak-to-strong, application [Easy2Hard (Sun et al., 2024b; Hase et al., 2024); Weak2Strong (Burns et al., 2023); Principle2Behavior (Bai et al., 2022b; Sun et al., 2023d) etc., application_work] ] ] ]
@@ 3 Aligning Through Inductive Bias
Currently, aligning through inductive bias is one of the most promising directions for achieving automated alignment. Inductive bias (Mitchell, 1980), are essentially assumptions or constraints that guide model learning and decision-making processes. By carefully selecting and implementing suitable inductive biases, we can steer models towards behaviors and decisions that are more likely to meet human standards and expectations, which can then generalize to unseen data distributions.
@@
After conducting a thorough review of the relevant literature, we find that current efforts towards self-improvement solely through language model itself can be decomposed into a set $\mathcal{H}$ of five inductive biases. These inductive biases fall into two broad categories: 1) Those stemming from inherent features of LLMs (§3.1), and 2) Those arising from the self-play dynamics of LLMs (§3.2).
For each type of inductive bias, we will begin by introducing its origin. Following this, we will enumerate the works that employ this inductive bias as a single-step policy improvement strategy. Next, we will discuss the works that iteratively train on these biases, aiming for continuous improvement. Finally, we will address the open research problems associated with the given inductive bias.
@@ 3.1 Inductive Bias from Features of LLMs
LLMs have intrinsic features that can function as inductive biases. These features primarily originate from the pre-training of deep Transformer networks on large datasets (H1, H3). Additionally, some biases result from initial alignment processes designed to improve the models’ helpfulness (H2). In this section, we summarize three key inductive biases, as depicted in Figure 3. It is important to note that these inductive biases are not entirely independent; instead, they represent three different perspectives on automated alignment.
LLMs possess intrinsic features that can act as inductive biases. These features largely arise from the pre-training of deep Transformer networks on massive datasets (H1, H3), while some also stem from preliminary alignment procedures aimed at enhancing the models’ helpfulness (H2). In this section, we will summarize three key inductive biases, as depicted in Figure 3. It is important to note that these inductive biases are not entirely independent; instead, they represent three different perspectives on automated alignment.
@@ 3.1.1 H1: Uncertainty as a indicator of helpfulness
Probability distributions from models can represent uncertainty. As Kadavath et al. (2022) discovered, when prompts are suitably designed, the responses obtained from pre-trained LLMs can be well-calibrated, and the degree of calibration can scale with the number of parameters and the number of exemplars. In other words, the higher the probabilities assigned by an LLM to a given answer, the more likely that answer is to be correct. This hypothesis has also been validated by Wang et al. (2021) and He et al. (2023). Similarly, Manakul et al. (2023) found a correlation between the probabilities output by aligned models and factuality.
In the machine learning literature, early applications of this inductive bias were evident in a series of works using self-training (Scudder, 1965) for semi-supervised learning (Nigam and Ghani, 2000; Amini and Gallinari, 2002). The basic paradigm of these works involves using learners trained on supervised data to continue learning from confidently classified unlabeled data, thereby enhancing supervised learning performance with unlabeled data. This approach has been applied in classification tasks with methods such as Pseudo-labeling (Lee et al., 2013; Ferreira et al., 2023) and Entropy Minimization (Grandvalet and Bengio, 2004). He et al. (2020) extended this approach to sequence generation NLP tasks, highlighting that biased sampling and noise perturbation are key factors for the success of self-training in these tasks. Pace et al. (2024) extended the paradigm of self-training to the alignment problem, improving the robustness of reward models by allowing them to iteratively learn from the highest-scoring and lowest-scoring answers in the candidate pool generated for a query.
@@
To perpetuate such enhancement, Self-Improve (Huang et al., 2023a) views CoT with Self-Consistency as a policy improvement operator, substantially improving the reasoning capabilities and potential of LLMs through iterative learning of the reasoning paths obtained from Self-Consistency. Zhang and Parkes (2023) demonstrate that LMs could self-improve on large number addition problems by curriculum “distilling” the CoT answers to the direct answers without explicit reasoning. Quiet-STAR (Zelikman et al., 2024) considers the advantage of the influence of rollout rationale on the probabilities of subsequent tokens as a feedback signal, encouraging the model to generate a more helpful implicit thought process using reinforcement learning techniques.
@@
Figure 4: Calibration plot observed by Kadavath et al. (2022); He et al. (2023); OpenAI (2023b); Zhang et al. (2024). The x-axis represents the probability associated with the model’s output, while the y-axis indicates the probability that the answer is correct. In the low probability region, the gray area may reflect the “don’t know” responses substituting some low-confidence answers. In the high probability region, the gray area signifies overconfidence. The comparison between pre-trained LLM and aligned LLM demonstrates that alignment for helpfulness can result in miscalibration, which is harmful to iterative self-improvement.
For aligned models, maintaining calibration and uncertainty remains crucial, as miscalibration can undermine the potential for iterative self-improvement. Numerous studies (Kadavath et al., 2022; He et al., 2023; OpenAI, 2023b; Zhang et al., 2024) have noted that the alignment process can impair the calibration of LLMs (as shown in Figure 4). This observation is reasonable for several reasons: 1) The current superficial alignment process aims to steer the model away from generating harmful or incorrect answers. This involves replacing the probability of incorrect answers with the probability of rejection responses to some extent, creating a gray area in the low probability region. 2) During the alignment process, the model also learns the response format. The increase in confidence regarding the response format can somewhat affect the confidence in the answers themselves (He et al., 2023). Moreover, the model relearns the correct answers, which can lead to overconfidence (represented by the gray area in the high probability region). The extremization of the probability distribution might be more pronounced in self-alignment, given that it is an iterative process involving self-sampling and training. This implies that all tokens are already in a very high probability distribution, making them more likely to be sampled as responses.
When the model becomes overconfident, it leads to a decrease in the diversity and exploratory nature of the model’s generated outputs. To mitigate this issue, one promising approach is to use inference-time interventions (e.g., high temperature (Kadavath et al., 2022), fidelity (Zhang et al., 2024)) to decrease the expected calibration error. Another potential solution is to filter the pseudo-labeled samples to avoid harmful repeated training, which requires understanding when unlabeled samples will be effective (Grandvalet and Bengio, 2004).
@@ 3.1.2 H2: LLMs can judge, critic, refine and more
Judge refers to determining the quality of model responses. These judgment standards are typically incorporated into instructions as principles or guidelines, enabling regulators to oversee LLM behavior in a more scalable manner (Bai et al., 2022b; Yuan et al., 2024) compared to heavy reliance on human annotators for feedback (Bai et al., 2022a). This approach allows for timely regulation, aiding in the flexible and controlled alignment process of language models, which can help prevent issues like reward hacking during iterative training (Sun et al., 2023c) and facilitate on-policy Reinforcement Learning Training (Guo et al., 2024b).
Self-judging can manifest in two primary forms: 1) Differentiating the relative quality of two responses (AI Feedback (Bai et al., 2022b)), resulting in evaluation outcomes represented as a partial order. For instance, Tan et al. (2023) employed prompts to compare which answer better adheres to the HHH principles. They then distilled the chosen option back into the model to further enhance its judging capability. Bai et al. (2022b) prompted the model to select superior responses based on sampled principles and subsequently employed a preference process to achieve a Pareto improvement of the model. 2) Providing an absolute score for a response (LLM-as-a-judge (Zheng et al., 2024b)), with the evaluation results in scalar form. Yao et al. (2023b), Besta et al. (2024) and Xie et al. (2023) introduced real-time evaluation modules for thought states during the reasoning process. These modules serve as a prior during the search process, assisting the model in exploring the action space for questions requiring deliberate thinking. Similarly, RAIN (Li et al., 2024g) utilizes a binary scoring prompt for self-evaluating whether the generation is likely to be harmful, thereby enhancing response safety combined with an inference-time tree search. Yuan et al. (2024) employed a five-point judge prompt to score the model’s instruction-response outputs, then converted the scores into a partial order for iterative training using DPO.
@@
Critique refers to generating modification suggestions. By leveraging LLM itself for critique, the suggestions can address errors and deficiencies, such as mistakes in summarization (Saunders et al., 2022) and machine translation (Fernandes et al., 2023), math reasoning (Lin et al., 2024b), and decision-making and programming tasks (Saunders et al., 2022). The suggestions can also pertain to abstract value criteria, such as the HHH related principles (Chen et al., 2024e; Bai et al., 2022b).
Refine refers to the ability that LLMs can improve the given text. Most of the work on self-refinement is based on natural language reasons provided by the critic module to modify the response (Bai et al., 2022b; Tan et al., 2023; Madaan et al., 2023). Some research also demonstrates the possibility of making modifications directly based on scalar rewards (Shinn et al., 2023). The less informative critic can be more challenging for LLMs since they must complete more information by themselves through reasoning. Another line of work uses LLMs to refine the prompts themselves (Fernando et al., 2023; Yang et al., 2023a).
Other: A helpful LLM can serve in various capacities to assist in the alignment process, effectively replacing human guidance. For example, it can vote on the quality of intermediate states of thought (Yao et al., 2023b), verify outcomes based on predicting conditions in questions (Weng et al., 2023), and automatically generate, filter (Yue et al., 2024b), and evolve instructions (Wang et al., 2023f; Li et al., 2024e; Xu et al., 2023a), among other tasks.
Regarding the persistence process, the improvements derived from these methods can be further distilled into the model through SFT (i.e., rejection sampling), DPO, RM-RL, and other techniques. Additionally, the judge / critique - refine process can be conducted iteratively.
@@
As the abilities of judging, critiquing, and refining are increasingly incorporated into model feedback and learning processes, there is a need for systematic evaluation studies of these capabilities. In this context, several research directions are worth pursuing:
@@
    Gathering empirical evidence to demonstrate that self critique-judge-refine abilities can enhance model performance in fair and reasonable experimental settings. Some works point out that the improvement may come from the use of stronger models (Sharma et al., 2024) and golden labels (Huang et al., 2023a).

### v2

@@ 1 Introduction
Recent years have witnessed the rapid advancements of large language models (LLMs), which have dramatically reshaped the landscape of artificial intelligence (Ouyang et al., 2022; Touvron et al., 2023; OpenAI, 2023c). Alignment is at the core of shaping behaviors of LLMs corresponding to human intentions and values (Yao et al., 2023a; Shen et al., 2023b), e.g., teaching LLMs to follow “helpful, harmless and honest (HHH)” principles during responding (Askell et al., 2021). As a result, increasing efforts have been made for aligning LLMs to meet the human requirements, which makes it a hotspot research direction in LLM era (Wang et al., 2023g, 2024g; Ji et al., 2023).
@@
To address these challenges, automated alignment has drawn great attention very recently (Yuan et al., 2024b; Chen et al., 2024g). Unlike previous methods that relied on human annotation to obtain alignment signals, the goal of automated alignment is constructing scalable and high-quality alignment systems with minimal human intervention. Therefore, automated alignment has the potential to address the core challenges posed by the rapid development of LLMs, where human annotation is either infeasible or extremely expensive. For automated alignment, the most crucial part is to find a scalable alignment signal that can replace human manually-created preference signals and remain effective amid the rapid development of LLMs.
@@
Bias §3, for tree=acquisition [From features of LLMs, [Uncertainty Filtering, acquisition [Self-Consistency (Wang et al., 2023e); Self-Improve (Huang et al., 2023a); West-of-N (Pace et al., 2024); etc., acquisition_work] ] [Self-Judge/Critic/Refine, acquisition [ Constitutional AI (Bai et al., 2022b); Tree-of-Thought (Yao et al., 2023b); Self-Rewarding (Yuan et al., 2024b); etc.,acquisition_work ] ] [Context Distillation, acquisition [Lab for Alignment (Askell et al., 2021); Dromedary (Sun et al., 2023d); Llama-2-Chat (Touvron et al., 2023); RLCD (Yang et al., 2024b); etc., acquisition_work] ] ] [From organization of LLMs, [Task Decomposition, acquisition [Least-to-Most (Zhou et al., 2023b); IDA (Christiano et al., 2018); etc., acquisition_work] ] [Self-play, acquisition [SPIN (Chen et al., 2024g); Consensus Game (Jacob et al., 2024); etc., acquisition_work] [Debate (Irving et al., 2018); SPAG (Cheng et al., 2024); etc., acquisition_work] ] ] ] [Behavior
Imitation §4, for tree=representation [ Instruction Construction, [ Unnatural Instructions (Honovich et al., 2023); Self-Instruct (Wang et al., 2023f); Evol-Instruct (Xu et al., 2024a); Humpback (Li et al., 2024f); etc., representation_work_2 ] ] [ Strong-to-Weak Distillation, [Response-guided, representation [ LLaMA-GPT4 (Peng et al., 2023); Stanford Alpaca (Taori et al., 2023); Ultrachat (Ding et al., 2023); etc., representation_work ] ] [Preference-guided, representation [ Zephyr (Tunstall et al., 2023); IterAlign (Chen et al., 2024e); Openchat (Wang et al., 2024b); etc., representation_work ] ] ] [ Weak-to-Strong Alignment, [ Weak2Strong (Burns et al., 2023); IaR (Somerstep et al., 2024); Liu and Alahi (2024); Hase et al. (2024); etc., representation_work_2 ] ] ] [Model
@@
Feedback §6, for tree=editing [ Social Interactions, editing [StableAlignment (Liu et al., 2023a); MoralDial (Sun et al., 2023a); SOTOPIA-$\pi$ (Wang et al., 2024f); etc., editing_work] ] [Human Shared-Values, editing [MGE (Klingefjord et al., 2024); Collective Constitutional AI (Anthropic, 2023) etc., editing_work] ] [Tool Execution, editing [Self-Debugging (Qiao et al., 2024); CodeRL (Qiao et al., 2024); SelfEvlove (Jiang et al., 2023); CRITIC (Gou et al., 2024); etc., editing_work ] ] [Embodied Environment, editing [GLAM (Carta et al., 2023); E2WM (Xiang et al., 2023); TWOSOME (Tan et al., 2024); etc., editing_work ] ] ] [Mechanism §7, for tree=application [Alignment
Mechanism, application [LIMA (Zhou et al., 2023a); Rethinking (Ren et al., 2024); URIAL (Lin et al., 2024a); ICL&IT (Duan et al., 2023); Behavior shift (Wu et al., 2023a) etc., application_work] ] [Inner Workings of Self-feedback,application [GV-consistency (Li et al., 2024g); CriticBench (Lin et al., 2024b); Self-Rewarding (Yuan et al., 2024b); Humback (Li et al., 2024f); LLM-as-a-Judge (Zheng et al., 2024b) etc., application_work] ] [Feasibility of Weak-to-strong, application [Easy2Hard (Sun et al., 2024b; Hase et al., 2024); Weak2Strong (Burns et al., 2023); Principle2Behavior (Bai et al., 2022b; Sun et al., 2023d) etc., application_work] ] ] ]
@@ 3 Aligning Through Inductive Bias
Currently, aligning through inductive bias is one of the most promising directions for achieving automated alignment. Inductive bias (Mitchell, 1980) is a set of essentially assumptions or constraints that guide model learning and decision-making processes. By carefully selecting and implementing suitable inductive biases, we can steer models towards behaviors and decisions that are more likely to meet human standards and expectations, which can then generalize to unseen data distributions.
@@
After conducting a thorough review of the relevant literature, we find that current efforts towards self-improvement solely through language model itself can be decomposed into a set $\mathcal{H}$ of five inductive biases. These inductive biases fall into two broad categories: 1) Those stemming from inherent features of LLMs (§3.1), and 2) Those arising from the organization of LLMs (§3.2).
For each type of inductive biases, we will begin by introducing its origin. Following this, we will enumerate the works that employ this inductive bias as a single-step policy improvement strategy. Next, we will discuss the works that iteratively train with it, aiming for continuous improvement. Finally, we will address open research problems associated with the given inductive bias.
@@ 3.1 Inductive Bias from Features of LLMs
LLMs possess intrinsic features that can act as inductive biases. These features largely arise from the pre-training of deep Transformer networks on massive datasets (H1, H3), while some also stem from preliminary alignment procedures aimed at enhancing the models’ helpfulness (H2). In this section, we will summarize three key inductive biases, as depicted in Figure 3. It is important to note that these inductive biases are not entirely independent; instead, they represent three different perspectives on automated alignment based on the features of LLMs.
@@ 3.1.1 H1: Uncertainty as a indicator of helpfulness
Probability distributions from models can represent uncertainty. As Kadavath et al. (2022) discovered, when prompts are suitably designed, the responses obtained from pre-trained LLMs can be well-calibrated, and the degree of calibration can scale with the number of parameters and the number of exemplars. In other words, the higher the probabilities assigned by an LLM to a given answer, the more likely that answer is to be correct. This hypothesis has also been validated by Wang et al. (2021) and He et al. (2023). Similarly, Manakul et al. (2023) found a correlation between the probabilities output by the aligned models and factuality.
In the machine learning literature, early applications of this inductive bias were evident in a series of works using self-training (Scudder, 1965) for semi-supervised learning (Nigam and Ghani, 2000; Amini and Gallinari, 2002). The basic paradigm of these works involves using learners trained on supervised data to continue learning from confidently classified unlabeled data, thereby enhancing supervised learning performance with unlabeled data. This approach has been applied in classification tasks with methods such as Pseudo-label (Lee et al., 2013; Ferreira et al., 2023) and Entropy Minimization (Grandvalet and Bengio, 2004). He et al. (2020) extended this approach to sequence generation NLP tasks, highlighting that biased sampling and noise perturbation are key factors for the success of self-training in these tasks. Pace et al. (2024) extended the paradigm of self-training to the alignment problem, improving the robustness of reward models by allowing them to iteratively learn from the highest-scoring and lowest-scoring answers in the candidate pool generated for a query.
@@
To solidify such enhancement, Self-Improve (Huang et al., 2023a) views CoT with Self-Consistency as a policy improvement operator, substantially improving the reasoning capabilities and potential of LLMs through iterative learning of the reasoning paths obtained from Self-Consistency. Zhang and Parkes (2023) demonstrate that LMs could self-improve on large number addition problems by curriculum “distilling” the CoT answers to the direct answers without explicit reasoning. Quiet-STAR (Zelikman et al., 2024) considers the advantage of the influence of rollout rationale on the probabilities of subsequent tokens as a feedback signal, encouraging the model to generate a more helpful implicit thought process using reinforcement learning techniques.
@@
Figure 4: Calibration plot observed by Kadavath et al. (2022); He et al. (2023); OpenAI (2023c); Zhang et al. (2024). The x-axis represents the probability associated with the model’s output, while the y-axis indicates the probability that the answer is correct. In the low probability region, the gray area may reflect the “don’t know” responses substituting some low-confidence answers. In the high probability region, the gray area signifies overconfidence. The comparison between pre-trained LLM and aligned LLM demonstrates that alignment for helpfulness can result in miscalibration, which is harmful to iterative self-improvement.
For aligned models, maintaining calibration and uncertainty remains crucial, as miscalibration can undermine the potential for iterative self-improvement. Numerous studies (Kadavath et al., 2022; He et al., 2023; OpenAI, 2023c; Zhang et al., 2024) have noted that the alignment process can impair the calibration of LLMs (as shown in Figure 4). This observation is reasonable for several reasons: 1) The current superficial alignment process aims to steer the model away from generating harmful or incorrect answers. This involves replacing the probability of incorrect answers with the probability of rejection responses to some extent, creating a gray area in the low probability region. 2) During the alignment process, the model also learns the response format. The increase in confidence regarding the response format can somewhat affect the confidence in the answers themselves (He et al., 2023). Moreover, the model relearns the correct answers, which can lead to overconfidence (represented by the gray area in the high probability region). The extremization of the probability distribution can be more pronounced in self-alignment (Wu et al., 2024b), given that it is an iterative process involving self-sampling and training. This implies that all tokens are already in a very high probability distribution, making them more likely to be sampled as responses.
When the model becomes overconfident, it leads to a decrease in the diversity and exploratory of the model’s generated outputs. To mitigate this issue, one promising approach is to use inference-time interventions (e.g., high temperature (Kadavath et al., 2022), fidelity (Zhang et al., 2024)) to decrease the expected calibration error. Another potential solution is to filter the pseudo-labeled samples to avoid harmful repeated training, which requires understanding when unlabeled samples will be effective (Grandvalet and Bengio, 2004).
@@ 3.1.2 H2: LLMs can judge, critic, refine and more
Judge refers to determining the quality of model responses. The judgment standards are typically incorporated into instructions as principles or guidelines, enabling regulators to oversee LLM behavior in a more scalable manner (Bai et al., 2022b; Yuan et al., 2024b) compared to heavy reliance on human annotators for feedback (Bai et al., 2022a). This approach allows for timely regulation, aiding in the flexible and controlled alignment process of language models, which can help prevent issues like reward hacking during iterative training (Sun et al., 2023c) and facilitate on-policy reinforcement learning training (Guo et al., 2024a).
Self-judging can manifest in two primary forms: 1) Differentiating the relative quality of two responses (AI Feedback (Bai et al., 2022b)), resulting in evaluation outcomes represented as a partial order. For instance, Tan et al. (2023) employ prompts to compare which answer better adheres to the HHH principles. They then distill the chosen option back into the model to further enhance its judging capability. Bai et al. (2022b) prompted the model to select superior responses based on sampled principles and subsequently employed a preference process to achieve a Pareto improvement of the model. 2) Providing an absolute score for a response (LLM-as-a-judge (Zheng et al., 2024b)), with the evaluation results in scalar form. Yao et al. (2023b), Besta et al. (2024) and Xie et al. (2023) introduce real-time evaluation modules for thought states during reasoning process. These modules serves as a prior during the search process, assisting the model in exploring the action space for questions requiring deliberate thinking. Similarly, RAIN (Li et al., 2024h) utilizes a binary scoring prompt for self-evaluating whether the generation is likely to be harmful, thereby enhancing response safety combined with an inference-time tree search. Yuan et al. (2024b) employ a five-point judge prompt to score the model’s instruction-response outputs, then converted the scores into a partial order for iterative training using DPO.
@@
Critique refers to generating modification suggestions. By leveraging LLM itself for critique, the suggestions can address errors and deficiencies, such as mistakes in summarization (Saunders et al., 2022), machine translation (Fernandes et al., 2023), math reasoning (Lin et al., 2024b), decision-making and programming tasks (Saunders et al., 2022). The suggestions can also pertain to abstract value criteria, such as the HHH related principles (Chen et al., 2024e; Bai et al., 2022b).
Refine refers to the ability that LLMs can improve the given text. Most of the work on self-refine is based on natural language reasons provided by the critic module to modify the response (Bai et al., 2022b; Tan et al., 2023; Madaan et al., 2023). Some research also demonstrates the possibility of making modifications directly based on scalar rewards (Shinn et al., 2023). The less informative critic can be more challenging for LLMs since they must complete more information by themselves through reasoning. Another line of work uses LLMs to refine the prompts themselves (Fernando et al., 2023; Yang et al., 2023a).
Other: A helpful LLM can serve in various capacities to assist in the alignment process, effectively replacing human guidance. For example, it can vote on the quality of intermediate states of thought (Yao et al., 2023b), verify outcomes based on predicting conditions in questions (Weng et al., 2023), and automatically generate, filter (Yue et al., 2024b), and evolve instructions (Wang et al., 2023f; Li et al., 2024f; Xu et al., 2023a), among other tasks.
Regarding the persistence process, the improvements derived from these methods can be further distilled into the model through SFT (i.e., Expert Iteration), DPO, RM-PPO, and other techniques. Additionally, the judge / critique - refine process can be conducted iteratively.
@@
As the abilities of judging, critiquing, and refining are increasingly incorporated into model feedback and learning processes, there is a need for systematic evaluation of these capabilities. In this context, several research directions are worth pursuing:
@@
    Gathering empirical evidence to demonstrate that self self-critique, judge, and refine abilities can enhance model performance in fair and reasonable experimental settings. Some works point out that the improvement may come from the use of stronger models (Sharma et al., 2024) and golden labels (Huang et al., 2023a).